A mass is oscillating on a spring with a period of 4.60 s. At t = 0 s the mass has zero speed and is at x = 8.30 cm. What is its speed at t = 2.50 s?
Answer #1

giaven that

T = 4.6

w = 2 pi / T = 2 pi / 4.6 = 1.365 rad/s

now speed at any instant

v^2 = w^2 ( A^2 - x^2)

v^2 = 1.365^2* ( 8.3^2 - 2.5^2)

v = 10.8 cm/s

==========

Comment in case any doubt, will reply for sure.. Goodluck